Coherent microwaves of wavelength 5.00cm enter a tall, narrow window in a building otherwise essentially opaque to the microwaves. If the window is 36.0cm wide, what is the distance from the central maximum to the first-order minimum along a wall 6.50m from the window?


1
Expert's answer
2021-03-02T07:07:28-0500

From the geometry of the problem (sketch of the problem), we can see that distance y from the central maximum to the first-order minimum to the first-order minimum can be found as

"tan\theta = \large\frac{y}{L}" (1)

where L is the distance from the window (slit) to the wall (screen) and angle "\theta" is the angle at which we are observing the minimum, which can be calculated using the expression.

"asin\theta = m\lambda" (2)



In equation (2) a is the width of the window, m = "\pm1, \pm2,..." is an order of the diffraction (in our case m = 1 because we are interested in the first-order minimum.) and "\lambda" is wavelength of the wave. Therefore, we can extract the angle "\theta" from (2).

"\theta = sin^{-1}\large(\frac{\lambda}{a})" (3)

and inert it to (1) to get y

"y = Ltan\theta" (Inserting (3))

"y = Ltansin^{-1}\large(\frac{\lambda}{a} )" = "6.5m*tansin^{-1}\large(\frac{5cm}{36cm})" = 0.912m = 91.2cm